OpenStudy (anonymous):

for a fixed t,the tangent line to a regular curve "alpha" at a point alpha(t) is the staight line u->alpha(t)+alpha'(t),where we delet the point of application of alpha'(t).find the tangent line to the helix alpha(t)=(2cost,2sint,t)at poin 0

OpenStudy (zehanz):

Do you mean at the point with t=0?

OpenStudy (anonymous):

yes

OpenStudy (zehanz):

So it is the point (2, 0, 0). alpha'(t)=(-2sint, 2cost, 1). What is alpha'(0)?

OpenStudy (anonymous):

( o,2,1) ?

OpenStudy (zehanz):

BTW, welcome to OpenStudy!! If you set t=0 in alpha'(t), you get (0, 2, 1) as tangent vector. You could use vectors to describe the tangent line in (2, 0, 0):\[\left(\begin{matrix}x \\ y \\ z\end{matrix}\right)=\left(\begin{matrix}2 \\ 0 \\0\end{matrix}\right)+\lambda \left(\begin{matrix}0 \\ 2\\ 1\end{matrix}\right)\]

OpenStudy (zehanz):

Or, you could use a system of equations: Because \(x=2,~ y=2+2\lambda, ~z=\lambda \), you could use this:\[x=2 , ~z=\frac{1}{2}y\]

OpenStudy (zehanz):

Sorry, of course, \(y=2\lambda\)

OpenStudy (anonymous):

thankyou zehanz so much and what is lemda?

OpenStudy (anonymous):

can u help me to find a unique curve?

OpenStudy (zehanz):

It is lambda, a Greek letter. It is used as a parameter: For every value of lambda, you get another point of the tangent line.